Find the equilibrium prices and quantities for Good X and Good Y by simultaneously solving given the following:

Qdx = 290 - 4Px + 2Py and Qsx = -20 + Px.

For Good Y: Qdy = 300 - 2Px - 6Py and Qsy = -15 + 2Py.
Answer #1

In equilibrium quantity demanded is equal to quantity supplied

For good x :

Qd = Qs

290 - 4Px + 2Py = - 20 + Px

5Px - 2Py = 310

For good y :

Qd= Qs

300- 2Px -6Py = - 15 + 2Py

2Px + 8 Py = 315

Solving the equations simultaneously for equilibrium price and quantity.

Px = 70.68

Py = 21.7

Similarly,

Qx= 50.68, round off to 51 units

Qy= 28.4, rounded off to 28 units

...